# Multi-modal atmospheric characterization of Pictoris b

### v1

@@ 4.2 Atmospheric structure
Exo-REM grids predict the equilibrium temperature profile and the mixing ratio profiles of the most important gases. Figure 4 shows the retrieved temperature profile for our best fit using the full model (in light blue) compared with the opacities of two cloud species, Fe and Mg2SiO4 (in gray and brown respectively), as a function of pressure. We observe a well-constrained pressure-temperature (P-T) profile, with a cloud deck forming around P bar, where the temperature becomes high enough for clouds to be optically thick and impact the observed spectrum. The implications in terms of variability will have to be explored with 3D simulations. Stellar irradiation is not accounted for in these models, as seen in the almost isothermal top of the atmosphere. However, for Pic b, a planet relatively close to its star (9.930.03 au), irradiation is likely to have an impact. Using predictions for Pic A (see Table 1 and Zwintz et al., 2019) for the stellar radius), we estimate the stellar irradiation received by Pic b to be F W.m-2, which is about twice the amount Jupiter receives today ( W.m-2). This increased irradiation could affect the planet’s atmospheric chemistry and, consequently, the retrieved parameters. The coupling between self-consistent atmospheric models and irradiation models is an area of atmospheric modeling that requires further exploration.
@@ 4.3 CO Isotopologue
Given the S/N of the GRAVITY spectrum () and the relatively high spectral resolution, we investigated whether we could retrieve some constraint on the isotopic ratio . With Exo-REM, we only manage to obtain a loose upper constraint of log(12C/13C) ¡ 1.08. Similarly, ATMO only constrained it up to log(12C/13C) ¡ 1.55. Since both grids does not explore the parameter space below log(12C/13C) = 1.0, it is hard to conclude on the exact upper bound, because we could just be seeing the end tail of a much larger posterior distribution. We used a similar strategy as Zhang et al. (2021) to assess the detection and its statistical significance. The statistical significance is directly obtain by comparing the Bayesian evidences, computing the Bayes factor:
for two runs using respectively the model with and without 13CO. The model without 13CO is the same model as the one exploring the 13CO abundance, but with a fixed log(12C/13C)=100. Focusing on Exo-REM, we obtained ; meaning that the observation favor a full model with 13CO at a significance level . To support this value, we computed the CCF between the observational residuals (i.e. ) and a 13CO model template (i.e. ). This CCF, performed around the 13CO band-heads (2.34-2.4 µm) compared with the re-normalized auto-correlation, is shown on Fig. 5. We see a tentative peak at 0 km.s-1 (S/N = 4.54) which could indicate a 13CO detection. However, the presence of strong residuals of the order of the variation induced by the 13CO makes this statement very dubious. Indeed, analyzing each GRAVITY observation separately revealed that the first night (2019/11/09) had significantly more telluric residuals. Inverting the data for each of the four nights individually provided only upper limits for the first night, while no constraints (i.e., a flat posterior) were obtained for the other three nights. We further investigate these detection limits and their significance using synthetic spectra in Appendix A.
@@ 4.4 Radial velocity
Radial velocities presented in Table 4 have been corrected from the Earth barycentric velocity. We note that Exo-REM’s and ATMO’s predictions are significantly shifted compared to those of BT-Settl and Sonora for the 2019/11/09 and 2022/01/24 epochs. Coincidentally, Exo-REM and ATMO are the only two grids with spectral sampling lower than GRAVITY beyond 2.3 µm, which may partially explain these discrepancies. Accounting for the systemic velocity of the host star ( km.s-1, Gontcharov, 2006) and propagating its error bars, we show on Fig. 7 the retrieved velocities for the BT-Settl model together with the orbital predictions from the astrometric measurements 555https://github.com/semaphoreP/whereistheplanet (Wang et al., 2021). They are relatively consistent with the predictions except the 2019/11/09 epoch.
@@ 5 Combining low and medium spectral resolution
This section focuses on the joint datasets described in Table 2, excluding the CRIRES data. Given the multi-modality of the datasets, two main approaches are followed for the inversion:
Free: we assume here that none of the datasets are perfectly calibrated in flux. ForMoSA applies a dilution factor ck to scale the model to the data flux, which is calculated analytically (Cushing et al., 2008) : c, where is the model, the data and their associated uncertainties.
Physically informed: In contrast to the approach above, we introduce constraints on the planet’s radius and use a new scaling parameter for each dataset . This parameter is a multiplicative factor defined as c (R/D). We impose a broad Gaussian prior on the radius ( ) extended from Chilcote et al. (2017), and use informed priors on parameters sensitive to flux calibration, such as log(g) = 4.180.13 dex from Lacour et al. (2021)’s dynamical mass measurements. To account for degeneracies, we further constrain with a Gaussian prior (), assuming the data are reasonably well-calibrated and allowing only small variations around 1.
We report in Tables 9, 10 and 11 the main characteristics of the posterior distributions when datasets are added iteratively. The main free parameters for each model are shown in Fig. 6. Both the Exo-REM and Sonora models remain statistically preferred over ATMO and BT-Settl, as supported by the Bayesian evidences and reduced (Table 6). We note that only the Exo-REM grid does not go off the edge of the grids when we add all the data, which in our view makes it the most robust in terms of derived physical parameters.
The best fit for each model are shown in Fig. 15. The overall pseudo-continuum shape is reproduced by all models but the residuals reveal noticeable differences. We interpret this as a consequence of the preponderance of some datasets containing the largest number of datapoints or with the largest S/N (e.g., GRAVITY and SINFONI) in the likelihood computation. As a consequence, datasets at lower resolutions exhibit stronger residuals, but which remain within the 1 envelope.
@@
Compared to the previous inversion on the GRAVITY datasets and considering mainly the Exo-REM and Sonora solutions, we note that:
@@
the radius was fitted separately using the ”Physically informed” runs using a relatively strong prior of 1.40.1 . Results are stored in the first column of Table 11. Exo-REM and Sonora end up finding more consistent radii of and respectively. The scaling parameters for the four models can be found on Fig. 9.
the metallicity remains super-solar when more datasets are added. For Exo-REM, we observe a shift with the inclusion of MATISSE data, where the metallicity decreases from to . Sonora shows no changes compared to the GRAVITY-only inversion, with at the grid border, suggesting very high metallicity.
@@
Figure 4 shows the evolution of the retrieved Exo-REM pressure-temperature profile when iteratively adding observations. We see that adding datasets has a noticeable effect on the profile, constraining significantly the 1 contour. The overall shape of the T/P profile is maintained. The two dominant cloud species (Fe and ) are overlaid as a function of pressure when using all datasets. They form in the photospheric region of the atmosphere.
for the detection of the CO isotopologue, when adding the SINFONI datasets to the GRAVITY one, we note that the Exo-REM models constrain a sub-solar log(12C/13C)=1.12, although a significant part of the tail of the distribution remains at the grid edge, as seen in Fig. 16. The Bayes factor between the model with and without 13CO is indicates that the model model with 13CO is preferred at . To support this, we also computed the CCF between the two residuals as done previously. The results are shown in Fig. 5 with the CCFs obtained using GRAVITY-alone, SINFONI-alone, and their combination. As in Sec. 4.3, we observe that the GRAVITY signal is heavily affected by residual telluric absorption. The SINFONI spectrum is less affected by this issue and the 13CO signal appears more pronounced. This weak and systematics-prone constraint does not support a robust 13CO detection at this stage.
finally, similar to Sect. 4.4, radial velocities are fitted separately for the four GRAVITY epochs and SINFONI. Results are stored in Table 11 in Appendix C. For GRAVITY, all models yield similar radial velocities, regardless of the added datasets, with values ranging between 5 km.s-1 and 15 km.s-1; in good agreement with predictions from the most recent orbital elements (Wang et al., 2021 and Fig. 7).
@@ 6.2 Results
For the combined inversion on the CRIRES+ data, we used the likelihood mapping from Ruffio et al. (2019) on the A and B nods, with the two nights processed separately to avoid re-interpolation. Each order was also fitted separately to avoid border issues when computing the continuum. We constrained the exploration of the radial velocity using the priors from Landman et al. (2024), with rv [20 km.s-1, 40 km.s-1]. For the M-band data, we avoided the B nod of the OB2 (second night) as well as the third order in all frames. All contribution where fitted in parallel for the K-band data while we only use the residual flux of the M-band data. We use the BT-Settl model since it is the only one capable of reaching both K- and M-band resolution requirement of 100 000 as shown on Fig. 1.
@@
Posteriors are provided in Table 7, and the corresponding best fit is shown in Fig. 8. In the first lower panel of Fig. 8, we observe that BT-Settl struggles to reproduce the the GPI YJ bands, a behavior noted earlier in Sect. 5 when fitting for the parameter. The model also shows deeper 12CO absorption lines, seen in both the GRAVITY and SINFONI data which is probably correlated to the high C/O retrieved. The last two lower panels display one of the spectral orders alongside the fitted model for the two CRIRES+ observations. The thermal contribution in the M-band is particularly pronounced, with the fitted planetary signal appearing almost flat, while in the K-band, we retrieve some broad planetary features.
@@
We do not observe any major changes when comparing the atmospheric parameters obtained using the ”Physically informed” approach from Sect. 5 with those retrieved here, even after adding the CRIRES+ data. This suggests that CRIRES+ does not provide any strong additional information on the atmospheric parameters. As expected, the absence of the planet’s continuum in the CRIRES+ datasets makes it difficult to constrain and log(g), and the values retrieved align with those predicted using only low- and medium-resolution datasets. However, changes in the posterior distributions of the C/O is more complex, as both the K-band (2.34–2.38 µm) and M-band (4.25–5.22 µm) feature strong CO absorption lines, which should influence the retrieved C/O. This is evident in the significantly lower C/O values obtained when inverting each CRIRES+ dataset individually, C/O = for the K-band and C/O = for the M-band (using priors of G(1700 K, 100 K) for and G(4.18 dex, 0.13 dex) for log(g)). In this simple framework we are fitting all possible orders which makes it likely that a significant amount of noise is introduced into the CCF signal used to compute the likelihood functions. This may reduce the prominence of the CO signal compared to that found in the GRAVITY, SINFONI, and MATISSE datasets. In particular in M-band, the large error-bars associated to the thermal noise render the likelihood function almost flat.
@@
The retrieved radial velocities for each CRIRES+ epoch agreed relatively well with orbital predictions although we do observe a significant spread around the expected value for the two epochs at K-band. Figure 7 compiles all the retrieved radial velocities for each epochs using BT-Settl. For the rotational velocity, we use the broadening function FastRotBroad from PyAstronomy666https://pyastronomy.readthedocs.io/en/latest/ jointly on the K- and M-bands CRIRES+ epochs. We obtained a value of v.sin(i) km.s-1, consistent at 1 with the values of Snellen et al. (2014) (253 km.s-1) and Kiefer et al. (2024) (25 km.s-1) and at 2 with the values of Landman et al. (2024) (19.91.0 km.s-1) and Parker et al. (2024) (222 km.s-1). One possible explanation for the discrepancy between the value retrieved here and those reported by Landman et al. (2024) and Parker et al. (2024) is that the forward model grids used in this study do not have a well-defined spectral resolution. Additional sources of uncertainty arise from the fact that, for computational efficiency, the grid is resampled and truncated in wavelength before the inversion. This inevitably affects both the retrieved radial and rotational velocities. Consequently, both the retrieved v.sin(i) should be interpreted as a broadening coefficient that accounts for these uncertainties and for the planet’s true rotational velocity.
@@ 7 Discussion
7.1 The multi-modal analysis: advantages and limitation
Using multiple datasets in a single and comprehensive framework is both powerful but also poses several challenges when wanting to infer on the physical properties of an exoplanet.
Photometry and low-resolution spectra primarily constrain and log(g), while high-resolution data lack the information on the emission continuum but capture the molecular features which are set by molecular abundances in the gas phase. When relying on a limited set of observations, the resulting posteriors can be degenerate. Combining multiple datasets within a joint likelihood framework can help break these degeneracies. This is illustrated in Fig. 17, where the bimodal solution obtained with Exo-REM (without 13CO) disappears once the MATISSE observation is included.
Known sources of error from each observation are properly propagated into the final posteriors; either through observational error bars or, when available, covariance matrices. Nevertheless, unaccounted noise and/or bias may still remain. This is clearly illustrated by the nonphysical ”uncertainties” retrieved on the atmospheric parameters in all our inversions. These uncertainties result from the propagation of observational errors (or covariances) but do not account for model systematics (Petrus et al., 2024). This issue is particularly evident in Figure 6, where the retrieved uncertainties are significantly smaller than the differences between models; even between the two best-fitting ones, Exo-REM and Sonora. Several studies (Czekala et al., 2015; Kawahara et al., 2022; de Regt et al., 2024, 2025; Rotman et al., 2025) have demonstrated the strength of Gaussian Processes in modeling unaccounted noise and spectral correlations, and in propagating them into the final posteriors, in the context of atmospheric modeling of exoplanets and brown-dwarfs. Thus, we argue that adapting these developments into our forward modeling framework is the next logical step in ForMoSA if we want to increase the robustness of our physical interpretations.

### v2

@@ 4.2 Atmospheric structure
Exo-REM grids predict the equilibrium temperature profile and the mixing ratio profiles of the most important gases. Figure 4 shows the retrieved temperature profile for our best fit using the full model (in light blue) compared with the opacities of two cloud species, Fe and Mg2SiO4 (in gray and brown respectively), as a function of pressure. We observe a well-constrained pressure-temperature (P-T) profile, with a cloud deck forming around P bar, where the temperature becomes high enough for clouds to be optically thick and impact the observed spectrum. The implications in terms of variability will have to be explored with 3D simulations. Stellar irradiation is not accounted for in these models, as seen in the almost isothermal top of the atmosphere. However, for Pic b, a planet relatively close to its star ( au), irradiation is likely to have an impact. Using predictions for Pic A (see Table 1 and Zwintz et al., 2019) for the stellar radius), we estimate the stellar irradiation received by Pic b to be F W.m-2, which is about twice the amount Jupiter receives today ( W.m-2). This increased irradiation could affect the planet’s atmospheric chemistry and, consequently, the retrieved parameters. The coupling between self-consistent atmospheric models and irradiation models is an area of atmospheric modeling that requires further exploration.
@@ 4.3 CO Isotopologue
Given the S/N of the GRAVITY spectrum () and the relatively high spectral resolution, we investigated whether we could retrieve some constraint on the isotopic ratio 12CO/13CO. With Exo-REM, we only manage to obtain a loose upper constraint of log(12C/13C) ¡ 1.08. Similarly, ATMO only constrained it up to log(12C/13C) ¡ 1.55. Since both grids does not explore the parameter space below log(12C/13C) = 1.0, it is hard to conclude on the exact upper bound, because we could just be seeing the end tail of a much larger posterior distribution. We used a similar strategy as Zhang et al. (2021) to assess the detection and its statistical significance. The statistical significance is directly obtain by comparing the Bayesian evidences, computing the Bayes factor,
for two runs using respectively the model with and without 13CO. The model without 13CO is the same model as the one exploring the 13CO abundance, but with a fixed log(12C/13C)=100. Focusing on Exo-REM, we obtained ; meaning that the observation favor a full model with 13CO at a significance level . To support this value, we computed the CCF between the observational residuals (i.e., data minus model without 13CO) and a 13CO model template (i.e., model with 13CO minus model without 13CO). This CCF, performed around the 13CO band heads (2.34–2.4 µm) compared with the re-normalized auto-correlation, is shown on Fig. 5. We see a tentative peak at 0 km.s-1 (S/N = 4.54) which could indicate a 13CO detection. However, the presence of strong residuals of the order of the variation induced by the 13CO makes this statement very dubious. Indeed, analyzing each GRAVITY observation separately revealed that the first night (2019/11/09) had significantly more telluric residuals. Inverting the data for each of the four nights individually provided only upper limits for the first night, while no constraints (i.e., a flat posterior) were obtained for the other three nights. We further investigate these detection limits and their significance using synthetic spectra in Appendix A.
@@ 4.4 Radial velocity
Radial velocities presented in Table 8 have been corrected from the Earth barycentric velocity. We note that Exo-REM’s and ATMO’s predictions are significantly shifted compared to those of BT-Settl and Sonora for the 2019/11/09 and 2022/01/24 epochs. Coincidentally, Exo-REM and ATMO are the only two grids with spectral sampling lower than GRAVITY beyond 2.3 µm, which may partially explain these discrepancies. Accounting for the systemic velocity of the host star ( km.s-1, Gontcharov, 2006) and propagating its error bars, we show on Fig. 7 the retrieved velocities for the BT-Settl model together with the orbital predictions from the astrometric measurements 111111https://github.com/semaphoreP/whereistheplanet (Wang et al., 2021). They are relatively consistent with the predictions except the 2019/11/09 epoch.
@@ 5 Combining low and medium spectral resolution
This section focuses on the joint datasets described in Table 2, excluding the CRIRES data. Given the multimodality of the datasets, two main approaches are followed for the inversion, described below.
Free: we assume here that none of the datasets are perfectly calibrated in flux. ForMoSA applies a dilution factor ck to scale the model to the data flux, which is calculated analytically (Cushing et al., 2008):
(4) where is the model, the data, and their associated uncertainties.
Physically informed: In contrast to the approach above, we introduce constraints on the planet’s radius, , and use a new scaling parameter, , for each dataset, . This parameter is a multiplicative factor defined as
(5) We imposed a broad Gaussian prior on the radius ( ) extended from Chilcote et al. (2017), and use informed priors on parameters sensitive to flux calibration, such as log(g) = dex from Lacour et al. (2021)’s dynamical mass measurements. To account for degeneracies, we further constrained with a Gaussian prior (), assuming the data are reasonably well-calibrated and allowing only small variations around 1.
We report in Tables 9, 10, and 14 the main characteristics of the posterior distributions when datasets are added iteratively. The main free parameters for each model are shown in Fig. 6. Both the Exo-REM and Sonora models remain statistically preferred over ATMO and BT-Settl, as supported by the Bayesian evidences and reduced (Table 12). We note that only the Exo-REM grid did not go off the edge of the grids when we added all the data, which, in our view, makes it the most robust in terms of derived physical parameters.
The best fits for each model are shown in Fig. 15. The overall pseudo-continuum shape is reproduced by all models but the residuals reveal noticeable differences. We interpret this as a consequence of the preponderance of some datasets containing the largest number of datapoints or with the largest S/N (e.g., GRAVITY and SINFONI) in the likelihood computation. As a consequence, datasets at lower resolutions exhibit stronger residuals, but which remain within the 1 envelope.
@@
Compared to the previous inversion on the GRAVITY datasets and considering mainly the Exo-REM and Sonora solutions, we note the following points.
@@
The radius was fitted separately using the ”physically informed” runs using a relatively strong prior of 1.4 0.1 . Results are stored in the first column of Table 14. Exo-REM and Sonora end up finding more consistent radii of and , respectively. The scaling parameters, , for the four models can be found in Fig. 9.
The metallicity remains super-solar when more datasets are added. For Exo-REM, we observe a shift with the inclusion of MATISSE data, where the metallicity decreases from to . Sonora shows no changes compared to the GRAVITY-only inversion, with [M/H] at the grid border, suggesting very high metallicity.
@@
Figure 4 shows the evolution of the retrieved Exo-REM pressure-temperature profile when we iteratively add observations. We see that adding datasets has a noticeable effect on the profile, constraining the 1 contour significantly. The overall shape of the pressure-temperature profile is maintained. The two dominant cloud species (Fe and Mg2SiO4) are overlaid as a function of pressure when using all datasets. They form in the photospheric region of the atmosphere.
For the detection of the CO isotopologue, when adding the SINFONI datasets to the GRAVITY one, we note that the Exo-REM models constrain a sub-solar log(12C/13C) , although a significant part of the tail of the distribution remains at the grid edge, as seen in Fig. 16. The Bayes factor between the model with and without 13CO is indicates that the model model with 13CO is preferred at . To support this, we also computed the CCF between the two residuals as done previously. The results are shown in Fig. 5 with the CCFs obtained using GRAVITY-alone, SINFONI-alone, and their combination. As in Sec. 4.3, we observe that the GRAVITY signal is heavily affected by residual telluric absorption. The SINFONI spectrum is less affected by this issue and the 13CO signal appears more pronounced. This weak and systematics-prone constraint does not support a robust 13CO detection at this stage.
Finally, similarly to what is described in Sect. 4.4, the radial velocities were fitted separately for the four GRAVITY epochs and SINFONI. Results are stored in Table 14 in Appendix B. For GRAVITY, all models yield similar radial velocities, regardless of the added datasets, with values ranging between 5 km.s-1 and 15 km.s-1; this is in good agreement with predictions from the most recent orbital elements (Wang et al., 2021 and Fig. 7).
@@ 6.2 Results
For the combined inversion on the CRIRES+ data, we used the likelihood mapping from Ruffio et al. (2019) on the A and B nods, with the two nights processed separately to avoid re-interpolation. Each order was also fitted separately to avoid border issues when computing the continuum. We constrained the exploration of the radial velocity using the priors from Landman et al. (2024), with rv [20 km.s-1, 40 km.s-1]. For the M-band data, we avoided the B nod of the OB2 (second night) as well as the third order in all frames. All contribution where fitted in parallel for the K-band data while we only use the residual flux of the M-band data. We use the BT-Settl model since it is the only one capable of reaching both K- and M-band resolution requirement of Rλ 100,000 as shown on Fig. 1.
@@
Posteriors are provided in Table 7, and the corresponding best fit is shown in Fig. 8. In the first lower panel of Fig. 8, we observe that BT-Settl struggles to reproduce the the GPI YJ bands, a behavior noted earlier in Sect. 5 when fitting for the parameter. The model also shows deeper 12CO absorption lines, seen in both the GRAVITY and SINFONI data which is probably correlated to the high C/O retrieved. The last two lower panels display one of the spectral orders alongside the fitted model for the two CRIRES+ observations. The thermal contribution in the M band is particularly pronounced, with the fitted planetary signal appearing almost flat, while in the K band, we retrieve some broad planetary features.
@@
We do not observe any major changes when comparing the atmospheric parameters obtained using the ”physically informed” approach from Sect. 5 with those retrieved here, even after adding the CRIRES+ data. This suggests that CRIRES+ does not provide any strong additional information on the atmospheric parameters. As expected, the absence of the planet’s continuum in the CRIRES+ datasets makes it difficult to constrain and log(g), and the values retrieved align with those predicted using only low- and medium-resolution datasets. However, changes in the posterior distributions of the C/O is more complex, as both the K band (2.34–2.38 µm) and M band (4.25–5.22 µm) feature strong CO absorption lines, which should influence the retrieved C/O. This is evident in the significantly lower C/O values obtained when inverting each CRIRES+ dataset individually, C/O = for the K band and C/O = for the M band (using priors of G(1700 K, 100 K) for and G(4.18 dex, 0.13 dex) for log(g)). In this simple framework we are fitting all possible orders which makes it likely that a significant amount of noise is introduced into the CCF signal used to compute the likelihood functions. This may reduce the prominence of the CO signal compared to that found in the GRAVITY, SINFONI, and MATISSE datasets. In particular in the M band, the large error-bars associated to the thermal noise render the likelihood function almost flat.
@@
The retrieved radial velocities for each CRIRES+ epoch agreed relatively well with orbital predictions although we do observe a significant spread around the expected value for the two epochs in the K band. Figure 7 compiles all the retrieved radial velocities for each epochs using BT-Settl. For the rotational velocity, we use the broadening function FastRotBroad from PyAstronomy131313https://pyastronomy.readthedocs.io/en/latest/ jointly on the K- and M-band CRIRES+ epochs. We obtained a value of km.s-1, consistent at 1 with the values of Snellen et al. (2014a) (25 3 km.s-1) and Kiefer et al. (2024) (25 km.s-1) and at 2 with the values of Landman et al. (2024) (19.9 1.0 km.s-1) and Parker et al. (2024) (22 2 km.s-1). One possible explanation for the discrepancy between the value retrieved here and those reported by Landman et al. (2024) and Parker et al. (2024) is that the forward model grids used in this study do not have a well-defined spectral resolution. Additional sources of uncertainty arise from the fact that, for computational efficiency, the grid is resampled and truncated in wavelength before the inversion. This inevitably affects both the retrieved radial and rotational velocities. Consequently, both the retrieved should be interpreted as a broadening coefficient that accounts for these uncertainties and for the planet’s true rotational velocity.
@@ 7 Discussion
7.1 The multimodal analysis: Advantages and limitations
Using multiple datasets in a single and comprehensive framework is both powerful, but also poses several challenges when wanting to infer on the physical properties of an exoplanet. Photometry and low-resolution spectra primarily constrain and log(g), while high-resolution data lack the information on the emission continuum but capture the molecular features which are set by molecular abundances in the gas phase. When relying on a limited set of observations, the resulting posteriors can be degenerate. Combining multiple datasets within a joint likelihood framework can help break these degeneracies. This is illustrated in Fig. 17, where the bimodal solution obtained with Exo-REM (without 13CO) disappears once the MATISSE observation is included.
Known sources of error from each observation are properly propagated into the final posteriors; either through observational error bars or, when available, covariance matrices. Nevertheless, unaccounted noise and/or bias may still remain. This is clearly illustrated by the nonphysical ”uncertainties” retrieved on the atmospheric parameters in all our inversions. These uncertainties result from the propagation of observational errors (or covariances) but do not account for model systematics (Petrus et al., 2024). This issue is particularly evident in Figure 6, where the retrieved uncertainties are significantly smaller than the differences between models; even between the two best-fitting ones, Exo-REM and Sonora.
Several studies (Czekala et al., 2015; Kawahara et al., 2022; de Regt et al., 2024, 2025; Rotman et al., 2025) have demonstrated the strength of Gaussian processes (GPs) in modeling unaccounted noise and spectral correlations, and in propagating them into the final posteriors, in the context of atmospheric modeling of exoplanets and brown dwarfs. Thus, we argue that adapting these developments into our forward modeling framework is the next logical step in ForMoSA if we want to increase the robustness of our physical interpretations.